# Enhanced strong interaction effect in synthetic spin-orbit coupling with mixed spin symmetry

### v2

@@ I Introduction
SOC in cold atoms is often discussed in the context of the mean-field regime, e.g. [8, 9, 10]. While the mean-field approximation has revealed interesting results, for example discovering the phase diagram [8], it imposes classical fields and ignores the quantum effects. To bridge single-particle physics and many-body physics in SOC systems, a mapping between cold atomic SOC systems and the Dicke model has been proposed [11]. However, the validity of the mapping is not obvious [12]. The assumption imposed for the mapping is that all particles occupy the same real space state, which automatically leads to the pseudo-spin state spanning symmetric spin space only. This assumption is commonly used for Bose gases with two internal degrees of freedom [13], however SOC couples pseudo-spin states and real space states and allows pseudo-spin states to get out of symmetric spin space. It is not clear how the lowest energy state is confined in symmetric spin space.
@@
In this work, we investigate the role of mixed spin symmetry in SOC systems. To this end, we consider the smallest system with mixed spin symmetry, a three-particle system with two internal states, and we focus on the comparison with two-particle systems. We have built the Hamiltonian of three bosons with two internal states in the presence of SOC by considering the second quantisation and have computed a few low-energy states. In comparison to two particle systems, we have found similarity in the appearance of mixed symmetric pseudo-spin states in the ground state, which is confirmed for strong interaction between anti-aligned pseudo-spins and reduced interaction energy. This is also seen in two-particle systems when the anti-symmetric pseudo-spin states emerge. Also, the emergence of non-symmetric pseudo-spin states challenges the mapping proposed in Ref. [11]. At the same time, we have also studied the spatial structure of the ground state by looking at the pair correlation and find that SOC assists contact interaction in inducing a strong interaction effect. Even for relatively small contact interactions, the pair correlation of the ground state has a structure similar to the strongly correlated Tonks-Girardeau gas. The pair correlation of the two-particle system is simply the total density distribution and does not display such a strong interaction effect.
@@ III Few interacting bosons with SOC
We present our formalism here. First, we introduce the Hamiltonian of our system in the first quantization basis and then rewrite it in the second quantization basis to induce the bosonic symmetry. This approach is often used to present the Hamiltonian of few particle systems, e.g. [15, 16, 17], and we refer interested readers to a detailed reference [18].
@@ III.1 Hamiltonian in the first quantisation basis
$\hat{H}_{0}$ $=\sum_{j=1}^{N_{p}}\left(\frac{\hat{p}_{j}^{2}}{2m}+\frac{m\omega^{2}\hat{x}_{j}^{2}}{2}+\frac{\hbar k}{m}\hat{p}_{j}\hat{\sigma}_{z}^{(j)}+\frac{\hbar\Omega}{2}\hat{\sigma}_{x}^{(j)}\right),$ (4)
where the third term is the SOC term and describes coupling between real space and pseudo-spin space and the fourth term is coherent coupling between pseudo-spins. The SOC term gives positive momentum to down spins and negative momentum to up spins. Without either of these terms, the single-particle Hamiltonian can be diagonalised with the basis of $\hat{\sigma}_{z}$ or $\hat{\sigma}_{x}$. The interaction Hamiltonian can be decomposed into pseudo-spin basis as
@@ III.2 Hamiltonian in the second quantization basis
$\epsilon_{i,j}$ $=\hbar\omega\left(n_{x}(i)+\frac{1}{2}\right)\delta_{i,j}+\frac{ik\xi}{\sqrt{2}}m_{\text{s}}(j)\delta_{m_{\text{s}}(i),m_{\text{s}}(j)}$
@@
where $\xi=\sqrt{\hbar/m\omega}$ is the trap length, $n_{x}(j)$ represents the label of eigenstate of the harmonic oscillator, and $m_{\text{s}}(j)$ represents the pseudo-spin state: down spins give $m_{\text{s}}(j)=-1$, and up spins give $m_{\text{s}}(j)=1$. The interaction part (5) is given by
@@ IV Numerical results
We construct the Hamiltonains (7)(III.2) and diagonalize them in a truncated space numerically. In this work, we fix $g_{\downarrow\downarrow}=g_{\uparrow\uparrow}=g$ to keep the symmetry between down and up spins. Moreover, to study competition between SOC and contact interaction, we turn on the anti-aligned pseudo-spin interaction $g_{\downarrow\uparrow}$ or the aligned pseudo-spin interaction $g$. The study of the two particle system has shown that these two cases give clear different properties in the ground state [14]. We use the trap energy $\hbar\omega$ as the energy unit and normalise the parameters with the trap energy $\hbar\omega$ and the trap length $\xi=\sqrt{\hbar/m\omega}$, and display normalised parameters in all the figures. Also, we fix $k\xi=4$ and set the cutoff of the truncated space as $M=50$. We discuss the justification of the cutoff $M$ in Appendix A, and our numerical code can be viewed in Ref. [19].
Figure 1: Energy difference $E_{j}-E_{0}$ between the first to fifth excited states and the ground state for (a) the non-interacting case, (b) $g_{\downarrow\uparrow}/\hbar\omega\xi=1$, and (c) $g_{\downarrow\uparrow}/\hbar\omega\xi=10$. The dashed line is the critical point $\Omega_{\text{c}}^{0}/\omega=2k^{2}\xi^{2}$ in free space, and the dotted line is when $E_{1}$ and $E_{0}$ start to deviate, $(E_{1}-E_{0})/\hbar\omega>0.01$.
@@ IV.1 Zero interactions $(g=g_{\downarrow\uparrow}=0)$
It is known that SOC systems can have three different ground state phases [8]: the stripe phase, the magnetised phase, and the single minimum phase. In the stripe phase, the ground state is a superposition of positive and negative momenta and thus displays an interference pattern. In the magnetised phase, the ground state acquires positive or negative momentum, and in the single minimum phase the spectrum only possess a single minimum, leading to zero momentum usually. Without contact interactions, only the stripe phase and the single minimum phase exist, and at the critical point some degeneracies are resolved. In free space the critical point between these two phases is given analytically by $\Omega_{\text{c}}^{0}/\omega=2k^{2}\xi^{2}$, and in trapped systems the density modulation modifies the critical value a lower value [20].
@@ IV.2 Anti-aligned interactions ($g_{\downarrow\uparrow}>0$ and $g=0$)
We consider non-zero positive anti-aligned interaction $g_{\downarrow\uparrow}>0$, i.e. when the unaligned pseudo-spin components suffer from repulsive interaction. As a result, in the absence of the coherent coupling only the aligned pseudo-spin states $\ket{\downarrow\downarrow\downarrow}$, $\ket{\uparrow\uparrow\uparrow}$ obtain the lowest energy. We have computed the energy spectrum for relatively weak interaction $g_{\downarrow\uparrow}/\hbar\omega\xi=1$ and for strong interaction $g_{\downarrow\uparrow}/\hbar\omega\xi=10$. In the former (latter) case, the energy shift due to the contact interaction is less (more) than $\hbar\omega$. It is found that the degeneracy between the ground state and the first excited state is resolved at a different value of $\Omega$ from the noninteracting case, and it depends on the interaction strength $g_{\downarrow\uparrow}$ (see the dotted lines in Fig. 1(b,c)). Such deviation from the noninteracting case is also seen in the two-particle system but slightly smaller than in the three-particle system. This is because in the three-particle system the contact interaction energy is larger than in the two-particle system, and the maximum energy shift due to contact interaction is $\hbar\omega$ for two particles and $3\hbar\omega$ for three particles. For $\Omega/\omega\simeq 40$, the energy spectrum for different $g_{\downarrow\uparrow}$ look similar to each other because the coherent coupling is so strong that the contact interaction contribution to the energy is negligible.
@@ IV.2.1 Pseudo-spin population
Furthermore, we have computed the population $p_{\text{M}1},p_{\text{M}2},p_{\text{M}3},p_{\text{M}4}$ of the mixed spin symmetric basis (II) (see the details of the calculation to Appendix B). Since these populations are the same due to our setting $g_{\downarrow\downarrow}=g_{\uparrow\uparrow}$, we have plotted the sum $p_{\text{M}}=p_{\text{M}1}+p_{\text{M}2}+p_{\text{M}3}+p_{\text{M}4}$ (see the dotted lines in Fig. 2(a,b)). We have found that a large amount of the mixed spin symmetric states appear in the ground state for the strong interaction case (b) in Fig. 2. Considering $p_{\text{unaligned}}=p_{\text{S}2+\text{S}3}+p_{\text{M}}$ with $p_{\text{S}2+\text{S}3}$ the population of |$\ket{S_{2}}$ and |$\ket{S_{3}}$, this leads to the non-monotonic growth of $p_{\text{aligned}}$, $p_{\text{unaligned}}$. This is also seen in the two-particle system, and in that case it is originated from the emergence of the anti-symmetric pseudo-spin states [14]. The interaction energy increases for increasing the contact interaction strength in general, but the anti-symmetric pseudo-spin states do not feel contact interactions. Thus, the appearance of the anti-symmetric pseudo-spin states for strong contact interactions suppresses the interaction energy. In three-particle systems, the mixed spin symmetric states is mixture of spin symmetric states and spin antisymmetric states, and this anti-symmetry reduces the interaction energy. We have calculated the interaction energy $E_{\text{int}}=\langle\hat{H}_{\text{int}}\rangle$ and plotted it as a function of $\Omega$, which displays a dent for strong interaction (see Fig. 2(c)). Finally, we have inspected the non-monotonic growth of $p_{\text{unaligned}}$ for a wide range of interaction $g_{\downarrow\uparrow}$ (see Fig. 3) and confirmed that the excess over $3/4$ is larger for stronger $g_{\downarrow\uparrow}$.
Figure 3: Excess of $p_{\text{unaligned}}$ over the population at the limit of $\Omega\to\infty$ as function of $\Omega/\omega$ and $g_{\downarrow\uparrow}/\hbar\omega\xi$, i.e. $p_{\text{unaligned}}-3/4$. Negative values are not shown for clear presentation.
@@ IV.2.2 Pair correlation
We study the pair correlations, and first let us focus on $g_{\downarrow\uparrow}/\hbar\omega\xi=1$ and change the coherent coupling strength $\Omega$ (see the middle panels at $g_{\downarrow\uparrow}=1$ in Fig. 4). For relatively small coherent coupling $\Omega/\omega=10$, the pair correlation is close to a Gaussian distribution, and the contact interaction affects the pseudo-spin populations but does not change the spatial structure of the ground state from the free particle case. By increasing the coherent coupling strength ($\Omega/\omega=20$), the shape is squished along $x=y$. At $\Omega/\omega=26$, a dent emerges at the centre ($x=y=0$). For increasing $\Omega$ more, a slit at $x=y$ becomes deeper, and two bumps remain. This slit is originated solely from the contact interaction, and the pair correlation remains in the same shape even for larger $\Omega$. It is consistent to the pseudo-spin populations $p_{\text{aligned}}$, $p_{\text{unaligned}}$, which reach plateau when $\Omega/\omega\simeq 30$.
To dig into the interplay between the contact interaction and the SOC, we fix the coherent coupling strength to $\Omega/\omega=26$ and change the contact interaction strength (see the panels at $\Omega=26$ in Fig. 4). By decreasing the contact interaction strength, the dent seen for $g_{\downarrow\uparrow}/\hbar\omega\xi=1$ becomes invisible. On the other hand, for increasing $g$ the pair correlation shows three bumps in $x>y$ and in $x<y$. This is a qualitatively similar feature to the TG gas other than the size of these bumps. It is interesting that such similarly to TG gas is seen even for relatively small interaction such as $g_{\downarrow\uparrow}/\hbar\omega\xi=3$. For larger $g_{\downarrow\uparrow}$, the same structure remains. Also, it is worth noting that there is no interaction between the same pseudo-spins. In the two-particle system, the counterpart of the pair correlation is density distribution $\rho_{2}(x_{1},x_{2})=|\Psi_{\text{GS}}(x_{1},x_{2})|^{2}$, and it was investigated in our previous work [14]. However, such strong interaction feature enhanced by SOC is not observed.
Figure 5: Energy difference $E_{j}-E_{0}$ between some excited states and the ground state for $g/\hbar\omega\xi=0.48$ (a) and $g/\hbar\omega\xi=2$ (b). The dashed line is the critical point $\Omega_{\text{c}}^{0}/\omega=2k^{2}\xi^{2}$ in free space, and the dotted line is when $E_{1}$ and $E_{0}$ start to deviate, $(E_{1}-E_{0})/\hbar\omega>0.01$.

### v3

@@ I Introduction
While SOC in cold atoms is often discussed in the context of the mean-field regime, e.g. [8, 9, 10], the regime beyond it has also been attracting attention, such as the regime of interactions that are strong enough to observe entanglement generation [11, 12, 13]. To bridge single-particle physics and many-body physics in SOC systems, a mapping between cold atomic SOC systems and the Dicke model has been proposed [14]. So far, the results have only been experimentally demonstrated in the mean-field regime. The work hinted that this mapping worked even beyond the mean-field regime, but the validity of the mapping is not obvious. The assumption imposed for the mapping is that all particles occupy the same real space state, which automatically leads to the pseudo-spin state spanning symmetric spin space only. This assumption is commonly used for Bose gases with two internal degrees of freedom [11], however SOC couples pseudo-spin states with real space states and allows pseudo-spin states to get out of symmetric spin space. Overlooking this contradiction, it was predicted by Refs. [15, 16] that spin squeezing is observed in the mean-field regime in the SOC system by applying this mapping. This essentially would mean that entanglement is generated in a non-interacting system, which is not physical.
@@
In this work, we investigate the role of mixed spin symmetry in SOC systems. To this end, we consider the smallest system with mixed spin symmetry, a three-particle system with two internal states, and we focus on the comparison with two-particle systems. We have built the Hamiltonian of three bosons with two internal states in the presence of SOC by considering the second quantisation and have computed a few low-energy states. In comparison to two particle systems, we have found similarity in the appearance of mixed symmetric pseudo-spin states in the ground state, which is confirmed for strong interaction between anti-aligned pseudo-spins and reduced interaction energy. This is also seen in two-particle systems when the anti-symmetric pseudo-spin states emerge. Also, the emergence of non-symmetric pseudo-spin states challenges the mapping proposed in Ref. [14]. At the same time, we have also studied the spatial structure of the ground state by looking at the pair correlation and find that SOC assists contact interaction in inducing a strong interaction effect. Even for relatively small contact interactions, the pair correlation of the ground state has a structure similar to the strongly correlated Tonks-Girardeau gas. The pair correlation of the two-particle system is simply the total density distribution and does not display such a strong interaction effect. It is worth noting that the strong interaction effect enhanced by SOC is analogous to Ref. [18], which shows that SOC amplifies the spin squeezing.
@@ III Few interacting bosons with SOC
We present our formalism here. First, we introduce the Hamiltonian of our system in the first quantization basis and then rewrite it in the second quantization basis to induce the bosonic symmetry. This approach is often used to present the Hamiltonian of few particle systems, e.g. [19, 20, 21], and we refer interested readers to a detailed reference [22]. Although we consider three particles in this work, the approaches explained here work for any particle number, and we keep the discussion general in this section.
@@ III.1 Hamiltonian in the first quantisation basis
$\hat{H}_{0}$ $=\sum_{j=1}^{N}\left(\frac{\hat{p}_{j}^{2}}{2m}+\frac{m\omega^{2}\hat{x}_{j}^{2}}{2}+\frac{\hbar k}{m}\hat{p}_{j}\hat{\sigma}_{z}^{(j)}+\frac{\hbar\Omega}{2}\hat{\sigma}_{x}^{(j)}\right),$ (4)
where $N$ is the particle number, $m$ is the mass of each particle, $\omega$ is the harmonic trap frequency, $k$ is the SOC strength, and $\Omega$ is the Raman coupling strength. The $\sigma_{x,z}^{(j)}$ are the Pauli matrices. The third term describes SOC between real space and pseudo-spin space, and the momentum of the $j$th particle is shifted as $\hat{p}_{j}\pm\hbar k$ for up spin and down spin, respectively. Experimentally, this coupling is induced by the momentum difference between the two Raman laser pulses [3], and $k$ is the projected wave number determined by the wavelength and the angle of intersection between the Raman lasers [23]. Without either of the third and the fourth terms, the single-particle Hamiltonian can be diagonalised with the basis of $\hat{\sigma}_{z}$ or $\hat{\sigma}_{x}$, although the Raman coupling cannot be turned off completely while keeping SOC due to the implementation of SOC. In this work, we consider repulsive contact interactions where particles repel only when two particles meet. The interaction Hamiltonian can be decomposed into three contributions as
@@
where $g_{\downarrow\downarrow},g_{\uparrow\uparrow},g_{\downarrow\uparrow}$ are the interaction strength.
@@ III.2 Hamiltonian in the second quantization basis
$\epsilon_{i,j}$ $=\hbar\omega\left(n_{x}(i)+\frac{1}{2}\right)\delta_{i,j}+\frac{ik\ell}{\sqrt{2}}m_{\text{s}}(j)\delta_{m_{\text{s}}(i),m_{\text{s}}(j)}$
@@
where $\ell=\sqrt{\hbar/m\omega}$ is the trap length, $n_{x}(j)$ represents the label of eigenstate of the harmonic oscillator, and $m_{\text{s}}(j)$ represents the pseudo-spin state: down spins give $m_{\text{s}}(j)=-1$, and up spins give $m_{\text{s}}(j)=1$. The interaction part (5) is given by
@@
While the mixed spin symmetry bases are defined in Eq. (II) in the first quantization representation, we adopt the second quantization representation for exact diagonalization. Therefore, we need to transform the second quantization representation of the Hamiltonians (7)(III.2) to the first quantization representation to see the spin symmetry. In the second quantization representation, we do not distinguish each particle and register only which energy level is occupied (see the left hand in Fig. 1). On the other hand, in the first quantization representation each particle is labelled, and we register which energy level each particle is in in order of the labels. Considering bosons, each basis in the second quantization representation is equivalent to superposition of all the combinations of the labelling (see the right hand in Fig. 1). For instance, replacing particles labelled by 1 and 2 does not change anything. In this way, the bases in the second quantization representation can be converted into the bases in the first quantization representation, which give us the quantum number and the pseudo-spin state of each basis, i.e., each basis is described as |$\ket{n_{1},\sigma_{1};n_{2},\sigma_{2};n_{3},\sigma_{3}}$ for $n_{1},n_{2},n_{3}=0,1,\ldots,M-1$ and $\sigma_{1},\sigma_{2},\sigma_{3}=\downarrow,\uparrow$. Finally, we transform the pseudo-spin basis $\downarrow$,$\uparrow$ , ↑ to the classified pseudo-spin basis $S_{j}$, $M_{j}$ by using the relations given in Eqs. (II)(II).
Figure 1: Example of transforming the second quantization representation to the first quantization representation. The left hand shows a state of three indistinguishable particles, and the right hand shows superposition between three states of three distinguishable particles symmetrized to fulfil the boson statistics.
@@ IV Numerical results
We construct the Hamiltonains (7)(III.2) for three particles and diagonalize them in a truncated space numerically. In this work, we fix $g_{\downarrow\downarrow}=g_{\uparrow\uparrow}=g$ to keep the symmetry between down and up spins. Moreover, to study competition between SOC and contact interaction, we turn on the anti-aligned pseudo-spin interaction $g_{\downarrow\uparrow}$ or the aligned pseudo-spin interaction $g$. The study of the two particle system has shown that these two cases give clear different properties in the ground state [17]. We use the trap energy $\hbar\omega$ as the energy unit and normalise the parameters with the trap energy $\hbar\omega$ and the trap length $\ell=\sqrt{\hbar/m\omega}$, for example the interaction strengths $g,g_{\downarrow\uparrow}$ are normalised with $\gamma\equiv\hbar\omega\ell$. Also, we fix $k\ell=4$ and set the cutoff of the truncated space as $M=50$. We discuss the justification of the cutoff $M$ in Appendix A, and our numerical code can be viewed in Ref. [24].
Figure 2: Energy difference $E_{j}-E_{0}$ between the first to fifth excited states and the ground state for (a) the non-interacting case, (b) $g_{\downarrow\uparrow}/\gamma=1$, and (c) $g_{\downarrow\uparrow}/\gamma=10$. The dashed line is the critical point $\Omega_{\text{c}}^{0}/\omega=2k^{2}\ell^{2}$ in free space, and the dotted line is when $E_{1}$ and $E_{0}$ start to deviate, $(E_{1}-E_{0})/\hbar\omega>0.01$.
@@ IV.1 Zero interactions $(g=g_{\downarrow\uparrow}=0)$
It is known that SOC systems can have three different ground state phases [8]: the stripe phase, the magnetised phase, and the single minimum phase. In the stripe phase, the ground state is a superposition of positive and negative momenta and thus displays an interference pattern. In the magnetised phase, the ground state acquires positive or negative momentum, and in the single minimum phase the spectrum only possess a single minimum, leading to zero momentum usually. Without contact interactions, only the stripe phase and the single minimum phase exist, and at the critical point some degeneracies are resolved. In free space the critical point between these two phases is given analytically by $\Omega_{\text{c}}^{0}/\omega=2k^{2}\ell^{2}$, and in trapped systems the density modulation modifies the critical value a lower value [25].
@@ IV.2 Anti-aligned interactions ($g_{\downarrow\uparrow}>0$ and $g=0$)
We consider non-zero positive anti-aligned interaction $g_{\downarrow\uparrow}>0$, i.e. when the unaligned pseudo-spin components suffer from repulsive interaction. As a result, in the absence of the coherent coupling only the aligned pseudo-spin states $\ket{\downarrow\downarrow\downarrow}$, $\ket{\uparrow\uparrow\uparrow}$ obtain the lowest energy. We have computed the energy spectrum for relatively weak interaction $g_{\downarrow\uparrow}/\gamma=1$ and for strong interaction $g_{\downarrow\uparrow}/\gamma=10$. In the former (latter) case, the energy shift due to the contact interaction is less (more) than $\hbar\omega$. It is found that the degeneracy between the ground state and the first excited state is resolved at a different value of $\Omega$ from the noninteracting case, and it depends on the interaction strength $g_{\downarrow\uparrow}$ (see the dotted lines in Fig. 2(b,c)). Such deviation from the noninteracting case is also seen in the two-particle system but slightly smaller than in the three-particle system. This is because in the three-particle system the contact interaction energy is larger than in the two-particle system, and the maximum energy shift due to contact interaction is $\hbar\omega$ for two particles and $3\hbar\omega$ for three particles. For $\Omega/\omega\simeq 40$, the energy spectrum for different $g_{\downarrow\uparrow}$ look similar to each other because the coherent coupling is so strong that the contact interaction contribution to the energy is negligible.
@@ IV.2.1 Pseudo-spin population
Furthermore, we have computed the population $p_{\text{M}1},p_{\text{M}2},p_{\text{M}3},p_{\text{M}4}$ of the mixed spin symmetric basis (II). Since these populations are the same due to our setting $g_{\downarrow\downarrow}=g_{\uparrow\uparrow}$, we have plotted the sum $p_{\text{M}}=p_{\text{M}1}+p_{\text{M}2}+p_{\text{M}3}+p_{\text{M}4}$ (see the dotted lines in Fig. 3(a,b)). We have found that a large amount of the mixed spin symmetric states appear in the ground state for the strong interaction case (b) in Fig. 3. Considering $p_{\text{unaligned}}=p_{\text{S}2+\text{S}3}+p_{\text{M}}$ with $p_{\text{S}2+\text{S}3}$ the population of |$\ket{S_{2}}$ and |$\ket{S_{3}}$, this leads to the non-monotonic growth of $p_{\text{aligned}}$, $p_{\text{unaligned}}$. This is also seen in the two-particle system, and in that case it is originated from the emergence of the anti-symmetric pseudo-spin states [17]. The interaction energy increases for increasing the contact interaction strength in general, but the anti-symmetric pseudo-spin states do not feel contact interactions. Thus, the appearance of the anti-symmetric pseudo-spin states for strong contact interactions suppresses the interaction energy. In three-particle systems, the mixed spin symmetric states is mixture of spin symmetric states and spin antisymmetric states, and this anti-symmetry reduces the interaction energy. We have calculated the interaction energy $E_{\text{int}}=\langle\hat{H}_{\text{int}}\rangle$ and plotted it as a function of $\Omega$, which displays a dent for strong interaction (see Fig. 3(c)). Finally, we have inspected the non-monotonic growth of $p_{\text{unaligned}}$ for a wide range of interaction $g_{\downarrow\uparrow}$ (see Fig. 4) and confirmed that the excess over $3/4$ is larger for stronger $g_{\downarrow\uparrow}$.
Figure 4: Excess of $p_{\text{unaligned}}$ over the population at the limit of $\Omega\to\infty$ as function of $\Omega/\omega$ and $g_{\downarrow\uparrow}/\gamma$, i.e. $p_{\text{unaligned}}-3/4$. Negative values are not shown for clear presentation.
@@ IV.2.2 Pair correlation
We study the pair correlations, and first let us focus on $g_{\downarrow\uparrow}/\gamma=1$ and change the coherent coupling strength $\Omega$ (see the middle panels at $g_{\downarrow\uparrow}=1$ in Fig. 5). For relatively small coherent coupling $\Omega/\omega=10$, the pair correlation is close to a Gaussian distribution, and the contact interaction affects the pseudo-spin populations but does not change the spatial structure of the ground state from the free particle case. By increasing the coherent coupling strength ($\Omega/\omega=20$), the shape is squished along $x=y$. At $\Omega/\omega=26$, a dent emerges at the centre ($x=y=0$). For increasing $\Omega$ more, a slit at $x=y$ becomes deeper, and two bumps remain. This slit is originated solely from the contact interaction, and the pair correlation remains in the same shape even for larger $\Omega$. It is consistent to the pseudo-spin populations $p_{\text{aligned}}$, $p_{\text{unaligned}}$, which reach plateau when $\Omega/\omega\simeq 30$.
To dig into the interplay between the contact interaction and the SOC, we fix the coherent coupling strength to $\Omega/\omega=26$ and change the contact interaction strength (see the panels at $\Omega=26$ in Fig. 5). By decreasing the contact interaction strength, the dent seen for $g_{\downarrow\uparrow}/\gamma=1$ becomes invisible. On the other hand, for increasing $g$ the pair correlation shows three bumps in $x>y$ and in $x<y$. This is a qualitatively similar feature to the TG gas other than the size of these bumps. It is interesting that such similarly to TG gas is seen even for relatively small interaction such as $g_{\downarrow\uparrow}/\gamma=3$. For larger $g_{\downarrow\uparrow}$, the same structure remains. Also, it is worth noting that there is no interaction between the same pseudo-spins. In the two-particle system, the counterpart of the pair correlation is density distribution $\rho_{2}(x_{1},x_{2})=|\Psi_{\text{GS}}(x_{1},x_{2})|^{2}$, and it was investigated in our previous work [17]. However, such strong interaction feature enhanced by SOC is not observed.
Figure 6: Energy difference $E_{j}-E_{0}$ between some excited states and the ground state for $g/\gamma=0.48$ (a) and $g/\gamma=2$ (b). The dashed line is the critical point $\Omega_{\text{c}}^{0}/\omega=2k^{2}\ell^{2}$ in free space, and the dotted line is when $E_{1}$ and $E_{0}$ start to deviate, $(E_{1}-E_{0})/\hbar\omega>0.01$.